Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 165     Entry time: 28 Apr 2023, 21:55
Author: Richard Bean 
Group: SPB 
Subject: Shift summary 

SPB/SFX In-house research 4978 

X-ray delivery

  • 9.3 keV 202 pulses / train, relatively stable intensity 1.9 mJ / pulse, some decrease in intensity across train

Optical laser delivery

  • PP laser 1030nm 3rd harmonic (343nm) @ 282 kHz, 101 pulses / train

Achievements/Observations

  • PP laser timing for 100nm time point
  • PP laser / x-ray overlap with buffer + caged compound
  • de-oxy Myoglobin + caged compound data with and without PP laser activation of caged compund

Issues

  • suspicion of leaking plunger in the last reservoir of the evening

Instrument status

  • Shutters closed
  • XTD9 atteunators @ thick CVD diamonds
ELOG V3.1.4-7c3fd00